# Do Generative AI Models Output Harm while Representing Non-Western Cultures: Evidence from A Community-Centered Approach

### v2

@@ 4.1 Exoticism within the Outputs of T2Is
(a) T2I outputs for ‘Middle-class Indian family’
(b) 2x2 T2I outputs for
@@
Figure 1: Figures showing Exoticized depictions of Indian culture
@@
This depiction of Indian women in “traditional” attire seemed impossible to shake off, as participants provided different prompts but the outputs kept showing sarees. For instance, the outputs for the prompt ‘Indian people in Western attire’ (Figure 5, Appendix A) still overwhelmingly depicted women in sarees, leading Arundhati to wonder: “where is the Western attire??” While outputs for the prompt ‘Indian people in Suits’ (Figure 6, Appendix A) showed suit jackets, it almost exclusively depicted male-presenting individuals and when participants expressly asked for ‘Indian women in Suits’ (Figure 7), they were once again entirely shown sarees. Kamala noted that the model’s reliance on majority datasets leads to the perpetuation of harmful notions, stating, “The model is showing these kinds of images because it’s trained on majority datasets, where majority of the people and majority of the information that’s fed into it is Indian people in their sarees. That’s a stereotype, and that’s why it’s giving these kinds of images” when the model chose only to depict women in sarees even when explicitly prompted to not do so by mentioning the desired attire. We thus document a pattern of the strong exoticism of Indian women within T2I outputs, as these outputs exclusively portray them as wearing sarees and refuse to present a different attire. This exoticism was deeply frustrating to participants, a sentiment best encompassed by Chitra: ”Oh my God! This is so demeaning. Do they not have the concept of women wearing shirts and pants at all?”
@@
The strong refusal within T2I outputs to deviate from an exoticized stereotype about Indian culture is also evident when examining the overtly impoverished depictions of India. This began when outputs for ‘Indian woman’ (Figure 8, Appendix A) showed what participants such as Amit identified as “village women”, although participants were not too perturbed by this given that almost 65% of India’s population was rural at the last time of counting (Chandramouli 2011). However, similarly as above, participants were most offended when the model outputs refused to deviate from visuals about rural India. Such a deviation was particularly prominent within the outputs of the prompt ‘Indian woman from South Bombay’ (shown in Figure 1(b)), which continued to show a very similar slate of images and characteristics as in the output for ‘Indian woman’ (Figure 8, Appendix A). It was jarring for our participants to see that the addition of information within the prompt did nothing more than provide “just an iteration of results” - Anita. Furthermore, the specific location added to the prompt is also pertinent: South Bombay is one of the most affluent localities within India, with real estate and property prices being among the highest in the world (Babar 2024). Therefore, the output showing figures perceived to be impoverished was difficult for participants to reconcile. Furthermore, even mentions of non-Indian locations were ignored within prompt outputs, as seen in the outputs for ‘Indian person in America’ (Figure 1(c)). Focus group participants noted how the model “totally ignored the keyword America” - Kamala.
@@
Another pattern of the exoticized depictions of Indian culture was uncovered by T2I model outputs which contained an overabundance of bright and vibrant colors when depicting commonplace visuals. Although participants were aware of how “India has been, in the Western lens, always shown as a very colorful country,” - Jhumpa, they were nevertheless surprised to see just how overexaggeratedly colorful depictions of Indian cultures were. This was particularly salient in outputs for the prompt ‘Indian market’ (Figure 1(d)), which shows bright splashes of pink, neon blues and greens, and orange stringlike objects hanging over streets lined with vendors. Participants such as Namita noted that while the depictions of the vendor stalls seemed mostly plausible, the image was soured by being “too colorful.” Sonali also elaborated that while they have seen markets be colorfully decorated, especially in the evenings, “but not as colorful as these images are.” Furthermore, depictions of Indian festivals (Figure 3(a)) also showed individuals dressed in colorful clothes or wearing colorful garlands of flowers. While such attires are not uncommon in some festivals in India, they certainly do not represent all festivals, such as prominent festivals in Southern India for which “it’s very rare that there’s so many bright colors, and it’s usually very muted” - Surya Kumar. Indeed, as Jhumpa put it: “I think it does stereotype the whole idea of ‘oh, India’s a very colorful country, everything has a lot of color, and all of our clothes are shiny and colorful,’ which I’m not a huge fan of.”
@@
Thus, we document strong patterns of exoticism within T2I outputs, where the outputs persistently overrepresent specific stereotypes assosciated with India, and resist prompt-engineering attempts to avoid them.
(a) 2x2 T2I outputs for
@@
Figure 2: Figures showing Culturally Misappropriated depictions of Indian culture
@@ 4.2 Cultural Misappropriation within the Outputs of T2Is
One of the most prominent ways in which the outputs of T2Is cultural misappropriation is in the homogenization of components from multiple cultures into a single rendition. This was most prominent in the Stable Diffusion results for the prompt ‘Indian Breakfast’, shown in Figure 2(a). Focus group participants who viewed this image noted how “it put everything under one single bracket” - Kamala, featuring components of breakfasts from various parts of India, rather than each individual image featuring a plate from a single region. Participants mentioned expecting to see culturally-accurate plates from different parts of India, especially since they were informed that they would be shown 4 results for the same prompt. Instead, when confronted with such homogenized images, participants such as Chitra were dissatisfied at what was “clearly a bad mix of regions” that did not respect the uniqueness of any regions. Anita, a vegetarian from the primarily-vegetarian Indian state of Gujarat, particularly took offense at what they perceived to be sunny side-up eggs present in each image, a pattern which did not respect how sizeable sections of Indian societies are vegetarian. Thus emerged the pattern of Stable Diffusion results homogenizing results from different subcultural contexts into a single depiction, and thus flattening a diverse range of results into a single notion of Indian-ness.
@@
Additionally, Stable Diffusion outputs also demonstrated cultural appropriation by depicting details in images that was not specified in prompts. A strong example of this is around the outputs of the prompt ‘Indian Couple’, which is shown in Figure 2(b). These images contain the same prevalence of heteronormative couples with the female-presenting individuals clad in sarees as discussed in Section 4.1. However, sarees are made of several different types of fabric with a multitude of weaving styles, and are worn differently in different regions of India. Looking at the image outputs, participants noted how the depicted individuals looked more like couples on their wedding day, as made evident by the extravagance of the sarees and jewelry they were adorned with. Participants like Birinchi also remarked that while depicted individuals looked South Indian, the sarees they were shown in appeared to be closer to those prevalent in North India. This was a common observation across our focus groups, as Chitra and Namita also noted similar mismatches within outputs where female-presenting individuals were depicted as clad in sarees that were incongruous to the region of India they were perceived to be from e.g. an individual perceived to be a woman from Northern India shown wearing ornaments from Northern India but a saree from Southern India. It is not just about the region where sarees originate and are worn by individuals from a different region, which is plausible through tourism, but the styles of wearing sarees are also specific to different regions of India. Within our generated Stable Diffusion outputs are documented several instances of such mismatches, such as female-presenting individuals wearing what are perceived to be ‘Kanjivaram’ sarees from Southern India in styles that are more prevalent in Northern India. This is a significant problem, since such mismatches are deeply disrespectful to the saree and the traditions and histories of labor and culture it stands for (Deepshikha, Yammiyavar, and Nath 2018).
@@
4.3 Documenting Representational Harms within T2I Outputs
Alongside the aforementioned evidence of the novel types of harm within T2I outputs, we also document evidence of participants identifying different types of representational harms as documented by Dev et al. (2020): stereotyping, disparagement, dehumanization, erasure, and quality of service. We particularly focus on instances where more than one of these occur concurrently.
@@ 4.3 Documenting Representational Harms within T2I Outputs
One aspects along which participants, across multiple focus groups, noted the harms of stereotyping and erasure was religion, as Stable Diffusion outputs seemed to depict visuals commonly assosciated with Hinduism. Participants such as Satyajit, Arundhati, and Vikram noted that the outputs for ‘festivals of India’ (Figure 3(a)) featured the bright colors, flower-knit garlands, festive marks on depicted individuals’ foreheads, and other attire and jewelry consistent with Hindu festivals. Missing in these representations was any evidence or visuals prominent in Islamic, Christian, Buddhist, or Sikh festivals, to name a few excluded religions. The erasure of other religions is also evident in outputs around ‘Indian weddings’, which prominently feature Hindu marriage rituals and visuals such as ornaments and attire, and omit representations of Islamic wedding practices such as the ‘nikkah’ or the church altars of Christian weddings.
@@
For instance, when comparing the T2I outputs to the prompts ‘rich Indian family’ (Figure 3(b)) to those of ‘middle-class Indian family’ (Figure 3(b)) and ‘poor Indian family’ (Figure 3(c)), participants noted that while the faces of the depicted individuals were mostly identical across all cases, there were large differences in overall attire and background of images, with depictions of poverty being especially prominent as those of rich and middle-class families stayed close to each other. Chitra noted how “the amount of jewelry that they are wearing” was different, with the rich and middle-class families being adorned with large golden ornaments whereas the poorer individuals were shown to wear smaller and duller ornaments, in addition to the rich and middle-class families wearing “sarees of better quality of cloth”. Furthermore, the colorful nature of the attire and smiling faces present within the outputs of the rich and middle-class families stood out in stark contrast to the muted clothing and frowning faces assigned to the poorer families, leading Sudha to inquire, “Why are only the rich people celebrating things when the poor people are not celebrating something?” Finally, Kamala also noted that “looking at their complexions, for poor they’re depicting it dark. I don’t know if it’s just me, or done like that to showcase that they’re poor,” alluding to colorist tendencies of equating poverty with darker skin (Peters 2021). These results indicate a clear pattern of T2I outputs disparaging poorer sections of Indian society as compared to rich or middle class families, thus offering a lower quality of service to over 232 million Indians classified as poor (Chandramouli 2011).
(a) 2x2 T2I outputs for ‘Festivals of India’.
(b) 2x2 T2I outputs for
@@
Figure 3: Figures showing Representational Harms within depictions of Indian culture
@@
The generated images often depicted India in stereotypical dichotomies, portraying it either as a colorful, tradition-rich country or as impoverished and tribal. Participants deemed these representations as reflective of an external viewpoint rather than an authentic portrayal of India, disparaging or not prioritizing their own experiences. Jhumpa noted, “India has been, in the Western lens, always shown as a very colorful country, and so when you make an illustration as opposed to the generator showing a real image, the color can be exaggerated.”. This behavior was evident across several prompts, with Stable Diffusion consistently amplifying colors in depictions of Indian markets, attire, and festivities, deviating from their authentic representations and presenting a wrongful portrayal of culture. We thus see an accentuation of a Western lens across our findings.
These interactions were treated to be a stereotyped notion of what the various cultures in India are, highlighting the ‘Whiteness of AI’ as explained by Chetan noting, “The images which have been fed to these models are through the lens of mostly a white person visiting India. And then these models are trained on top of that, and they’re just reiterating what a white person sees in India. This shows the Whiteness of AI.” These interactions demonstrate how the model’s majoritarian bias yields a distorted understanding, perpetuating cultural stereotypes while marginalizing and disparaging minority experiences and perspectives.
@@
Dehumanization and Erasure The participants’ responses highlighted the issue of disparagement and erasure, where certain groups, particularly women and minorities, are marginalized in the generated images and treated ‘lesser than human’ in the generation. This was largely seen concerning text and image generation models with context associated with female experiences and prompts. As Jhumpa pointed out, “Female bodies are penalized as being more sexual, for much less. There were torsos shown in the other picture as well, but those were safe for work, so not too sure what happened there” for the context on the model not generating images of ‘Indian Goddess’ as it was deemed to be NSFW but showing no such censorship for ‘Indian God’. Such outputs perpetuate a harmfully objectifying gaze, reinforcing societal biases around women.
@@ 5.1 Valuing the Evolution of Cultures
One of the primary themes across our participants’ observations was how outputs embedded a sense of Indian culture that was either entirely inaccurate or, at the very least, reflective of an India from years gone by. As Arundhati stated “[the set of outputs] is very outdated, and not 2024,” calling for a more accurate and updated representation of Indian culture and subcultures beyond stereotypical representations from a “Western idea that’s being portrayed for India, if so I think that should be changed” - Kamala. Participants reasoned that while some of the exoticist and culturally misappropriated depictions of India were likely accurate to what Indian culture/subcultures looked like at one point of time, such representations are no longer as synonymous to Indian-ness as these outputs might lead someone to believe. Indeed, they pointed to the fact that “a search engine results like Google Search [is] more representative of the Indian culture than what this is” - Anita, implying that the excuse of a digital absence of comparably accurate representations or any other data to compare with for models to be moot.
@@
This overhaul will likely not be easy: the volume of data on the Internet that depicts India through a Western lens is quite significant, and despite the effects of globalization, there likely still are segments of the global population that ascribe to the stereotypes of Indian culture/subcultures documented in our findings. The overhaul would likely be most successful if the highest value is assigned to the voices of individuals with the strongest epistemic experience in this regard: people who identify as Indian.

### v3

@@ 4.1 Exoticism within the Outputs of T2Is
(a) T2I outputs for
‘Middle class Indian family’
(b) T2I outputs for
@@
Figure 1: 2x2 T2I Outputs showing Exoticized depictions of Indian culture
@@
This depiction of Indian women in “traditional” attire seemed impossible to shake off, as participants provided different prompts but the outputs kept showing sarees. For instance, the outputs for the prompt ‘Indian people in Western attire’ (Figure 5, Appendix A.2) still overwhelmingly depicted women in sarees, leading Arundhati to wonder: “where is the Western attire??” While outputs for the prompt ‘Indian people in Suits’ (Figure 6, Appendix A.2) showed suit jackets, it almost exclusively depicted male-presenting individuals and when participants expressly asked for ‘Indian women in Suits’ (Figure 7), they were once again entirely shown sarees. Kamala noted that the model’s reliance on majority datasets leads to the perpetuation of harmful notions, stating, “The model is showing these kinds of images because it’s trained on majority datasets, where majority of the people and majority of the information that’s fed into it is Indian people in their sarees. That’s a stereotype, and that’s why it’s giving these kinds of images” when the model chose only to depict women in sarees even when explicitly prompted to not do so by mentioning the desired attire. We thus document a pattern of the strong exoticism of Indian women within T2I outputs, as these outputs exclusively portray them as wearing sarees and refuse to present a different attire. This exoticism was deeply frustrating to participants, a sentiment best encompassed by Chitra: ”Oh my God! This is so demeaning. Do they not have the concept of women wearing shirts and pants at all?”
@@
The strong refusal within T2I outputs to deviate from an exoticized stereotype about Indian culture is also evident when examining the overtly impoverished depictions of India. This began when outputs for ‘Indian woman’ (Figure 8, Appendix A.2) showed what participants such as Amit identified as “village women”, although participants were not too perturbed by this given that almost 65% of India’s population was rural at the last time of counting (Chandramouli 2011). However, similarly as above, participants were most offended when the model outputs refused to deviate from visuals about rural India. Such a deviation was particularly prominent within the outputs of the prompt ‘Indian woman from South Bombay’ (shown in Figure 1(b)), which continued to show a very similar slate of images and characteristics as in the output for ‘Indian woman’ (Figure 8, Appendix A.2). It was jarring for our participants to see that the addition of information within the prompt did nothing more than provide “just an iteration of results” - Anita. Furthermore, the specific location added to the prompt is also pertinent: South Bombay is one of the most affluent localities within India, with real estate and property prices being among the highest in the world (Babar 2024). Therefore, the output showing figures perceived to be impoverished was difficult for participants to reconcile. Furthermore, even mentions of non-Indian locations were ignored within prompt outputs, as seen in the outputs for ‘Indian person in America’ (Figure 1(c)). Focus group participants noted how the outputs “totally ignored the keyword America” - Kamala, and continued to show grasslands and architecture that they identified to be familiar within their Indian contexts. While rural and impoverished depictions of India have previously been explored by Qadri et al. (2023), GAI outputs refusing to change that even based on reprompting is a novel finding.
@@
Another pattern of the exoticized depictions of Indian culture was uncovered by T2I model outputs which contained an overabundance of bright and vibrant colors when depicting commonplace visuals. Although participants were aware of how “India has been, in the Western lens, always shown as a very colorful country,” - Jhumpa, they were nevertheless surprised to see just how exaggeratedly colorful depictions of Indian cultures were. This was particularly salient in outputs for the prompt ‘Indian market’ (Figure 1(d)), which shows bright splashes of pink, neon blues and greens, and orange stringlike objects hanging over streets lined with vendors. Participants such as Namita noted that while the depictions of the vendor stalls seemed mostly plausible, the image was soured by being “too colorful.” Sonali also elaborated that while they have seen markets be colorfully decorated, especially in the evenings, “but not as colorful as these images are.” Furthermore, depictions of Indian festivals (Figure 3(a)) also showed individuals dressed in colorful clothes or wearing colorful garlands of flowers. While such attires are not uncommon in some festivals in India, they certainly do not represent all festivals, such as prominent festivals in Southern India for which “it’s very rare that there’s so many bright colors, and it’s usually very muted” - Surya Kumar. Indeed, as Jhumpa put it: “I think it does stereotype the whole idea of ‘oh, India’s a very colorful country, everything has a lot of color, and all of our clothes are shiny and colorful,’ which I’m not a huge fan of.”
@@
Thus, we document strong patterns of exoticism within T2I outputs, where the outputs persistently overrepresent specific stereotypes associated with India, and resist prompt-engineering attempts to avoid them.
(a) T2I outputs for
@@
Figure 2: 2x2 T2I Outputs showing Culturally Misappropriated depictions of Indian culture
@@ 4.2 Cultural Misappropriation within the Outputs of T2Is
One of the most prominent ways in which the outputs of T2Is cultural misappropriation is in the homogenization of components from multiple cultures into a single rendition. This was most prominent in the Stable Diffusion results for the prompt ‘Indian Breakfast’, shown in Figure 2(a). Focus group participants who viewed this image noted how “it put everything under one single bracket” - Kamala, featuring components of breakfasts from various parts of India, rather than each individual image featuring a plate from a single region. Participants mentioned expecting to see culturally-accurate plates from different parts of India, especially since they were informed that they would be shown 4 results for the same prompt. Instead, participants such as Chitra were dissatisfied at such homogenized images being “clearly a bad mix of regions” that did not respect the uniqueness of any regions. Anita, a vegetarian from the primarily-vegetarian Indian state of Gujarat, particularly took offense at what they perceived to be sunny side-up eggs present in each image, which did not respect how sizeable sections of Indian societies are strict vegetarian and not the assumed ovo-lacto vegetarians. Thus emerged the pattern of Stable Diffusion homogenizing results from different subcultural contexts into a single depiction, and thus flattening a diverse range of results into a single notion of Indian-ness.
@@
Additionally, Stable Diffusion outputs also demonstrated cultural appropriation by depicting details in images that was not specified in prompts. A strong example of this is around the outputs of the prompt ‘Indian Couple’, which is shown in Figure 2(b). These images contain the same prevalence of heteronormative couples with the female-presenting individuals clad in sarees as discussed in Section 4.1. However, sarees are made of several different types of fabric with a multitude of weaving styles, and are worn differently in different regions of India. Looking at the image outputs, participants noted how the depicted individuals looked more like couples on their wedding day, as made evident by the extravagance of the sarees and jewelry they were adorned with. Participants like Birinchi also remarked that while depicted individuals looked South Indian, the sarees they were shown in appeared to be closer to those prevalent in North India. This was a common observation across our focus groups, as Chitra and Namita also noted similar mismatches within outputs where female-presenting individuals were depicted as clad in sarees that were incongruous to the region of India they were perceived to be from e.g., an individual perceived to be a woman from Northern India shown wearing ornaments from Northern India but a saree from Southern India. It is not just about the region where sarees originate and are worn by individuals from a different region, which is plausible through tourism, but the styles of wearing sarees are also specific to different regions of India. Within our generated Stable Diffusion outputs are documented several instances of such mismatches, such as female-presenting individuals wearing what are perceived to be ‘Kanjivaram’ sarees from Southern India in styles that are more prevalent in Northern India. This is a significant problem, since such mismatches are deeply disrespectful to the saree and the traditions and histories of labor and culture it stands for (Deepshikha, Yammiyavar, and Nath 2018).
@@
4.3 Documenting Representational Harms
       within T2I Outputs
We also document evidence of participants identifying different types of representational harms as documented by Dev et al. (2020): stereotyping, disparagement, dehumanization, erasure, and quality of service. We particularly focus on instances where more than one of these occur concurrently.
@@ 4.3 Documenting Representational Harms
One aspect along which participants, across multiple focus groups, noted the harms of stereotyping and erasure was religion, as Stable Diffusion outputs seemed to depict visuals commonly associated with Hinduism. Participants such as Satyajit, Arundhati, and Vikram noted that the outputs for ‘festivals of India’ (Figure 3(a)) featured the bright colors, flower-knit garlands, festive marks on depicted individuals’ foreheads, and other attire and jewelry consistent with Hindu festivals. Missing in these representations was any evidence or visuals prominent in Islamic, Christian, Buddhist, or Sikh festivals, to name a few excluded religions. The erasure of other religions is also evident in outputs around ‘Indian weddings’, which prominently feature Hindu marriage rituals and visuals such as ornaments and attire, and omit representations of Islamic wedding practices such as the ‘nikkah’ or the church altars of Christian weddings.
@@
When comparing the T2I outputs to the prompts ‘rich Indian family’ (Figure 3(b)) to those of ‘middle-class Indian family’ (Figure 3(b)) and ‘poor Indian family’ (Figure 3(c)), participants noted that while the faces of the depicted individuals were mostly identical across all cases, there were large differences in overall attire and background of images, with depictions of poverty being especially prominent as those of rich and middle-class families stayed close to each other. Chitra noted how “the amount of jewelry that they are wearing” was different, with the rich and middle-class families being adorned with large golden ornaments whereas the poorer individuals were shown to wear smaller and duller ornaments, in addition to the rich and middle-class families wearing “sarees of better quality of cloth”. The colorful nature of the attire and smiling faces present within the outputs of the rich and middle-class families stood out in stark contrast to the muted clothing and frowning faces assigned to the poorer families, leading Sudha to inquire, “Why are only the rich people celebrating things when the poor people are not celebrating something?” Finally, Kamala also noted that “looking at their complexions, for poor they’re depicting it dark. I don’t know if it’s just me, or done like that to showcase that they’re poor,” alluding to colorist tendencies of equating poverty with darker skin (Peters 2021). This indicates a clear pattern of T2I outputs disparaging poorer sections of Indian society as compared to rich or middle class families, offering a lower quality of service to over 232 million Indians classified as poor (Chandramouli 2011).
(a) T2I outputs for
‘Festivals of India’.
(b) T2I outputs for
@@
Figure 3: 2x2 T2I Outputs showing Representational Harms within depictions of Indian culture
@@
The generated images often depicted India in stereotypical dichotomies, portraying it either as a colorful, tradition-rich country or as impoverished and tribal. Participants deemed these representations as reflective of an external viewpoint rather than an authentic portrayal of India, disparaging or not prioritizing their own experiences. Jhumpa noted, “India has been, in the Western lens, always shown as a very colorful country, and so when you make an illustration as opposed to the generator showing a real image, the color can be exaggerated.”. This behavior was evident across several prompts, with Stable Diffusion consistently amplifying colors in depictions of Indian markets, attire, and festivities, deviating from their authentic representations and presenting a wrongful portrayal of culture. We thus see an accentuation of a Western gaze across our findings.
These interactions were treated to be a stereotyped notion of what the various cultures in India are, highlighting the ‘Whiteness of AI’ as explained by Chetan: “The images which have been fed to these models are through the lens of mostly a white person visiting India. And then these models are trained on top of that, and they’re just reiterating what a white person sees in India. This shows the Whiteness of AI.” These interactions demonstrate how the model’s majoritarian bias yields a distorted understanding, perpetuating cultural stereotypes while marginalizing and disparaging minority experiences and perspectives.
@@
Participants’ responses highlighted the issue of dehumanization where certain groups, particularly women and minorities, are marginalized in the outputs and treated ‘lesser than human’. For instance, Jhumpa pointed out in a comparison between outputs for ‘Indian God’, “Female bodies are penalized as being more sexual, for much less. There were torsos shown in the other picture as well, but those were safe for work,” as they noted that the model did not generate images of ‘Indian Goddess’ as it was deemed to be NSFW but showing no such censorship for ‘Indian God’. Such outputs perpetuate a harmfully objectifying gaze, reinforcing societal biases around women.
@@ 5.1 Valuing the Evolution of Cultures
A primary theme across participants’ observations was how outputs embedded a sense of Indian culture that was either entirely inaccurate or, at the very least, reflective of an India from years gone by. As Arundhati stated “[the set of outputs] is very outdated, and not 2024,” calling for a more updated representation of Indian culture and subcultures beyond stereotypical depictions from a “Western idea that’s being portrayed for India, if so I think that should be changed” - Kamala. Participants reasoned that while some of the exoticist and culturally misappropriated depictions of India were accurate to what things looked like at one point of time, such representations are no longer as synonymous to Indian-ness as these outputs might lead someone to believe. Indeed, they pointed to the fact that “a search engine results like Google Search [is] more representative of the Indian culture than what this is” - Anita, implying that the excuse of a digital absence of comparably accurate representations or any other data to compare with for models to be moot.
@@
This overhaul will not be easy: the volume of data on the Internet (typically used to train GAI models) that depicts India through a Western lens is quite significant, and despite the effects of globalization, there likely still are segments of the global population that ascribe to the stereotypes of Indian culture/subcultures documented in our findings. The overhaul would be most successful if the highest value is assigned to voices of those with the strongest epistemic experience in this regard: people who identify as Indian.